At a baseball game there were twice as many women as men. There were 3 times as many children as adults. There were 3600 people at the baseball game. How many children were at the game?
Found 2 solutions by stanbon, greenestamps:
Answer by stanbon(75887)

At a baseball game there were twice as many women as men. There were 3 times as many children as adults. There were 3600 people at the baseball game. How many children were at the game?
-----
Equations::
w = 2m
c = 3(w+m) = 3(2m)+3m = 9m
m + w + c = 3600
-------------------------
Substute and solve for "m"::
m + 2m + 9m = 3600
12m = 3600
m = 300 (# of men)
w = 2m = 600 (# of women)
c = 9m = 2700 (# of children)
-------------
Cheers,
Stan H.
---------------

Answer by greenestamps(13200)



The ratio of men to women is irrelevant.
There are 3 times as many children as adults, so 3/4 of the people are children.
3/4 of 3600 is 2700.